OpenStudy (anonymous):

A plane flying with a constant speed of 150 km/h passes over a ground radar station at an altitude of 3 km and climbs at an angle of 30°. At what rate is the distance from the plane to the radar station increasing a minute later? (Round your answer to the nearest whole number.) what I suppose to differentiate?

OpenStudy (anonymous):

Find X = the rate of change of distance =speed cos 30 = 150 / X X = 150 / cos 30 Once you find X divide it by 60 to get the rate of change a minute later